OpenStudy (anonymous):

Identify the 29th term of the arithmetic sequence 6, 14, 22 … how do i do this

OpenStudy (anonymous):

Find the pattern in the sequence. Each term increases by 8. 6+8=14, 14+8=22. To find the nth term of the sequence, take (n-1)*8 and add it to the first term. Example: the third term can be calculated by (3-1)*8+6, or 16+6, which is 22, which you can see from the question is in fact the third term.

OpenStudy (anonymous):

what is n and add what to what term???

OpenStudy (anonymous):

i think the answer is 230 but idk how to get that

OpenStudy (anonymous):

Do you recognize the pattern in the sequence? How each term is 8 more than the term before it? You just want to figure out what happens if you do that 28 times (to get the 29th term).

OpenStudy (anonymous):

yah but i dnt understand the equation of doing that

OpenStudy (anonymous):

If you want to take 6 and add 8 to it a given number of times, how would you do that? For example, if you want to take 6 and add 8 three times, 6+8+8+8, what's an easier way to do that? You can take 6+(3*8) to add all three eights at once. 6+8+8+8=6+(3*8)=30. Now just apply that on a larger scale.